# Real-Time Wayfinding Assistant for Blind and Low-Vision Users

### v1

@@ 2.4. Gaps in Existing Research
Despite major advances in assisted navigation and AI-powered accessibility technologies, issues such as processing efficiency, real-time flexibility, usability, and scalability continue to impede their practical adoption by blind and low-vision (BLV) people.
One significant barrier is the processing cost of AI-powered systems. Many depth estimation and object identification methods need a lot of computing power, which makes them unsuitable for mobile deployment. While some technologies solve these restrictions, relying on specialist hardware like depth cameras and LiDAR raises costs and reduces accessibility (See et al., 2022). Although lightweight AI models have been proposed, they frequently forfeit accuracy, limiting their practical application.
Adaptability to changing environments is another key concern. Many navigation systems rely on pre-mapped layouts or presume static surroundings, rendering them ineffective in real-world urban environments with ever-changing impediments. Efforts to integrate external infrastructure, such as public surveillance cameras, aim to address this issue, but reliance on fixed installations limits scalability (Jain et al., 2023). Similarly, inertial-based indoor navigation systems exhibit localization drift over time, resulting in positioning errors (Tsai et al., 2024).
Excessive or incorrect information frequently causes confusion and reduces trust in assistive tools (Gonzalez Penuela et al., 2024). AI hallucinations, where generated responses are misleading or overly detailed, present additional concerns, particularly in multimodal accessibility tools (Seo et al., 2024). These inconsistencies diminish the reliability of AI-based assistance, making it difficult for users to rely on such systems in critical situations.
Customization and user flexibility are frequently ignored. Numerous assistive solutions adopt a uniform approach, neglecting the varied needs of users. Although multimodal interfaces have enhanced accessibility for data visualization, they often fall short in providing customization for various degrees of visual impairment (Zhang et al., 2024). Furthermore, studies emphasize the differing information preferences of people who are blind from birth compared to those who lost their sight later, but many navigation technologies do not accommodate these differences (Hoogsteen et al., 2022).
Scalability and inclusivity continue to be significant issues. Educational materials frequently fall short in providing adequate support for BLV learners, necessitating additional assistive tools for fundamental accessibility (Shaheen, 2024). The lack of universally applicable solutions worsens exclusion, highlighting the necessity for more inclusive and scalable technologies.
@@ 3.1. Navigation Schemes
To guide BLV individuals effectively, we consider multiple navigation schemes that define movement direction and obstacle-free pathfinding:
@@ 3.1.3. Clock Hour Hand (DOF 13)
Inspired by the clock-face representation, this method divides the navigation space into 13 distinct directions, mimicking how blind individuals often conceptualize orientation (e.g., ”move toward 12 o’clock” for straight ahead). This approach balances accuracy and user intuitiveness, making it a practical solution for real-world applications.
Figure 1. Diagram of three different navigation schemes to define movement direction.
[…]
A woman and a girl in white dresses sit in an open car.
Figure 2. Flow diagram of all our approaches, including the proposed one, for identifying the free path for BLV.
[…]
A woman and a girl in white dresses sit in an open car.
@@ 3.2. AI-Assisted Navigation Approaches
This study explores six different navigation approaches, five of which leverage AI models for obstacle avoidance and pathfinding. In contrast, the proposed Approach 6 introduces a computationally efficient image processing algorithm using Depth-First Search (DFS) to determine the longest free path. These approaches vary in input data, processing techniques, and model complexity, aiming to balance accuracy, efficiency, and real-time usability.
@@ 3.2.2. Approach 2: Depth Images to VLMs
64×64 depth images are used instead of RGB images to enhance depth perception. Depth information allows VLMs to understand spatial relationships more effectively, improving obstacle detection and navigation decisions. VLMs still require significant computational power for real-time processing, making this approach less efficient for low-resource devices.
@@ 3.2.4. Approach 4: Patch-Based Depth Representation to LLMs
Instead of using full-resolution depth images, this method processes depth data in a 5×5 grid of patch values, significantly reducing input size while retaining essential spatial details. The processed depth patches are fed into LLMs, which analyze the segmented depth information for navigation. Although this method lowers computational requirements, the reduction in resolution may impact accuracy in complex environments.
@@ 3.2.5. Approach 5: Classification-Based Pathfinding using Vision Transformers (ViT)
This approach formulates pathfinding as a classification task, where a Vision Transformer (ViT) processes images to predict the best navigation direction. ViTs are powerful deep-learning models that capture long-range dependencies within an image. To enable this, we manually annotated a dataset with directional labels and trained a ViT model to learn the mapping between image features and navigation decisions. However, ViTs require extensive computational resources and large-scale training data, limiting their feasibility for real-time deployment in assistive navigation applications.
3.3. Approach 6: DFS-Based Pathfinding from Patched Depth Images (Proposed)
@@
To address these challenges, we propose a computationally efficient image processing algorithm for pathfinding. Our method leverages depth images obtained from monocular depth estimation and utilizes a 15×15 patch-based depth representation. Instead of relying on deep learning, we implement a Depth-First Search (DFS) algorithm to traverse the depth image and identify the longest free path, ensuring a clear and safe route for navigation. As illustrated in Figure 4, this approach offers several advantages, including lightweight computation, real-time performance, and greater interpretability.
@@
DFS-Based Optimal Pathfinding Algorithm:
@@
DEFINE recursive DFS to explore paths:
@@
IF avg_intensity $<$ threshold OR no valid moves THEN
@@
IF top rows reached OR intensity diff. $>$ threshold THEN
@@
RECURSE DFS with updated path
@@
CALL DFS from (start_row, start_col)
@@
Figure 3. Diagram for DFS-based pathfinding algorithm illustrating path exploration and optimal route selection.
@@ 4. Results and Discussion
Approach 4 (Depth(5x5(patch_value)) $\rightarrow$ LLM) Phi-3.5-mini-instruct 47.90 18.33 3.257 40.00 32.38 3.257
@@
Approach 6 (Depth(15x15(patch_value)) $\rightarrow$ DFS)(Proposed) Pathfinder Algorithm 39.81 32.04 0.377 25.78 37.67 0.399
@@
Approach 4 (Depth(5x5(patch_value)) $\rightarrow$ LLM) Phi-3.5-mini-instruct 40.50 15.33 3.72 42.00 13.33 3.72
@@
Approach 6 (Depth(15x15(patch_value)) $\rightarrow$ DFS)(Proposed) Pathfinder Algorithm 36.45 34.67 1.136 26.00 38.33 1.332
@@ 4.1. Evaluation Metrics
To evaluate the performance of the proposed directional prediction algorithm, we consider three key metrics: Mean Absolute Error (MAE), Accuracy, and Average Response Time.
@@ 4.1.1. Mean Absolute Error (MAE)
MAE measures the average deviation between the predicted navigational direction and the ground truth, both expressed in clock-hour format. The ground truth directions were manually annotated separately for 300 images from the NYU Depth V2 (Nathan Silberman and Fergus, 2012) dataset (indoor scenes) and 300 images from the DIODE (Vasiljevic et al., 2019) dataset (outdoor scenes). For example, if the ground truth direction is annotated as 11:00 (i.e., $330^{\circ}$), and the predicted direction is 10:00 (i.e., $300^{\circ}$), then the angular deviation is $|330^{\circ}-300^{\circ}|=30^{\circ}$, which corresponds to a 1-hour error. Formally, MAE is calculated as:
@@ 4.1.3. Average Response Time
Figure 4. An in-depth visualization of our proposed method for identifying the optimal path using the Pathfinder algorithm, applied to a sample input from the DIODE (Vasiljevic et al., 2019) dataset.
[…]
A woman and a girl in white dresses sit in an open car.
@@ 4.2. Indoor Scene Performance
The performance comparison for indoor scenes, as shown in Table 2, reveals key insights into the effectiveness and efficiency of the proposed and baseline approaches. Notably, the proposed Approach 6 (Depth(15x15(patch_value)) $\rightarrow$ DFS) demonstrates competitive performance across key metrics, balancing mean absolute error (MAE), accuracy, and average response time.
For the indoor scene with 30 test images, Approach 6 achieves an MAE of 25.78 degrees with an accuracy of 37.67% and a notably fast average response time of 0.399 seconds. While the MAE value is slightly better than some LLM-based approaches (e.g., Approach 4 (Depth(5x5(patch_value)) $\rightarrow$ LLM) with 37.5 degrees MAE using gemini-1.5-pro), the proposed method significantly outperforms others in terms of computational efficiency. The proposed approach achieves an average response time of 0.399 seconds, which is significantly faster than VLM and LLM-based methods, where response times range between 3.094 and 5.75 seconds for other approaches. This indicates that the DFS-based Pathfinder Algorithm is highly optimized for rapid processing, making it a suitable choice for real-time navigation applications.
Moreover, the classification-based Approach 5 (ViT-Base) exhibits the lowest response time of 0.087 seconds, but its MAE of 52.5 degrees and accuracy of 13.29% indicate a trade-off between speed and prediction quality. This suggests that while ViT-based approaches are computationally efficient, they may lack the precision required for fine-grained obstacle navigation tasks.
@@ 4.3. Outdoor Scene Performance
Table 3 presents the performance comparison for outdoor scenes, where the proposed Approach 6 again demonstrates remarkable efficiency and improved accuracy under specific conditions. For the set of 30 test images, Approach 6 achieves an MAE of 26.00 degrees with an accuracy of 38.33%, outperforming all other approaches in both metrics. This represents a significant improvement over the best-performing VLM-based approach (Approach 4 with gemini-1.5-pro), which achieved an MAE of 41.00 degrees and an accuracy of 16.67%.
@@
The efficiency of the proposed approach is further highlighted by its response time of 1.332 seconds, which, while slower than the ViT-based method (0.091 seconds), is still markedly faster than all VLM and LLM-based methods. This suggests that the Pathfinder Algorithm strikes an effective balance between accuracy and computational efficiency, making it well-suited for real-world applications where rapid decision-making is critical.
@@
Interestingly, the proposed Approach 6 also performs consistently across indoor and outdoor environments, with only a slight increase in response time when transitioning to outdoor scenes. This consistency indicates the robustness of the DFS-based method in diverse environments, which is essential for navigation systems assisting blind and low-vision (BLV) individuals.

### v2

@@ 2.4. Gaps in Existing Research
Despite substantial advances in assistive navigation and AI-powered accessibility technology, major gaps remain, limiting their real-world impact for blind and low-vision (BLV) people.
First, several approaches encounter computational and hardware limitations. Depth estimation and object identification algorithms frequently need significant computing resources, rendering them unsuitable for reliable internet-free deployment. Moreover, these systems frequently rely on specialized hardware such as depth cameras or LiDAR, which raises costs and limits accessibility (See et al., 2022).
Second, many systems struggle with environmental flexibility. Navigation solutions that rely on pre-mapped patterns or static environments fail in dynamic urban settings where barriers are continuously shifting. Infrastructure-based solutions, such as those employing public cameras (Jain et al., 2023a), show potential but confront issues with scalability and large-scale dependability. Similarly, inertial-based indoor navigation systems are susceptible to time drift, diminishing long-term accuracy (Tsai et al., 2024).
Third, concerns about trust and information quality persist. BLV users frequently become frustrated when systems deliver excessive or misleading information. AI hallucinations, which provide false or overly detailed responses, might exacerbate these challenges and diminish trust in key situations (Gonzalez Penuela et al., 2024; Seo et al., 2024).
Fourth, personalization and user flexibility are frequently undeveloped. Many technologies use a one-size-fits-all concept, ignoring variations in blindness onset, mobility tactics, and individual preferences. Multimodal interfaces (e.g., touch, audio, haptics) have improved accessibility for data visualization (Zhang et al., 2024), but they still fail to meet various user needs. Studies reveal that early-blind and late-blind users generally require different types of environmental information, yet most systems fail to account for these distinctions (Hoogsteen et al., 2022).
Finally, the larger challenges of scalability and inclusion remain unaddressed. Educational environments, for example, continue to underserve BLV learners, with students frequently learning assistive technology literacy outside formal teaching (Shaheen, 2024). This absence of universally adaptable and inclusive solutions contributes to ongoing marginalization and emphasizes the importance of designs that are adaptable across contexts, from schools to busy urban streets.
These limitations suggest that future research should look beyond improving technological accuracy and instead address concerns of trust, personalization, social context, and equity. These gaps are both technological and sociotechnical, emphasizing the importance of efficiency, adaptability, usability, and inclusive system design. This entails designing navigation and accessibility systems not only as technological artifacts, but also as components of a larger interconnected ecosystem in which usability, collaboration, and inclusivity are equally as vital as algorithmic performance.
@@ 3.1. Navigation Schemes
To guide BLV individuals effectively, we consider three navigation schemes that define movement direction and obstacle-free pathfinding:
@@ 3.1.3. Clock Hour Hand (DOF 13)
Inspired by the clock-face representation, this method divides the navigation space into 13 distinct directions, mimicking how blind individuals often conceptualize orientation (e.g., ”move toward 12 o’clock” for straight ahead). This approach balances accuracy and user intuitiveness.
Figure 1. Visualization of three navigation schemes to define navigation direction.:
Figure 2. Block diagram of all our approaches for identifying the best path for BLV navigation.:
In this study, explore six different navigation approaches, five of which leverage AI models for obstacle avoidance and pathfinding. In contrast, the proposed Approach 6 introduces a computationally efficient image-processing algorithm that employs a novel pathfinding technique to determine the longest free path. These approaches differ in terms processing methods, and model complexity, with the goal of balancing accuracy, efficiency, and real-time usability.
@@ 3.2.2. Approach 2: Depth Images to VLMs
64×64 depth images are used instead of RGB images to enhance depth perception. Depth information allows VLMs to understand spatial relationships more effectively, improving obstacle detection and navigation decisions. VLMs still require significant computational power for real-time processing, making this approach less efficient for low-resource settings.
@@ 3.2.4. Approach 4: Patch-Based Depth Representation to LLMs
Instead of using full-resolution depth images, this method processes depth data in a 15×15 grid of patch values, significantly reducing input size while retaining essential spatial details. The processed depth patches are fed into LLMs, which analyze the segmented depth information for navigation. Although this method lowers computational requirements, the reduction in resolution may impact accuracy in complex environments.
@@ 3.2.5. Approach 5: Classification-Based Pathfinding using Vision Transformers (ViT)
In this approach, pathfinding is formulated as a classification problem, where the system predicts the best navigation direction from an input image. To achieve this, we developed a custom dataset labeled with clock-hand directions, representing the free space available in each scene. A visual annotation tool was designed that overlays a virtual clock hand spanning from the 9 o’clock to the 3 o’clock positions on the image. Annotators selected the direction of free space that aligned most closely with the hour-hand, and this was recorded as the ground-truth navigation direction (Figure 3).
Figure 3. Direction annotation using a clock-hand overlay, where the nearest free space to the virtual hour hand is labeled as the ground-truth navigation direction.
Two annotators manually labeled each image using a web-based interface, and their annotations were cross-verified using Cohen’s kappa agreement procedure (Vieira et al., 2010) by three more reviwers. They further inspected the dataset and validated the final labels. The resulting dataset comprised 270 training images and 30 testing images. Figure 4 depicts the end-to-end annotation process used to generate the dataset labels.
Figure 4. End-to-end annotation workflow used to generate the dataset labels.
We then employed a supervised learning strategy, fine-tuning the Vision Transformer (ViT-Base) model on this annotated dataset. The model was trained to learn the mapping between extracted image features and the labeled navigation directions. While ViTs are effective at capturing long-range dependencies within images, their requirement for significant computational resources and large-scale data makes them less practical for real-time deployment in assistive navigation contexts. Nonetheless, this experiment demonstrates the feasibility of formulating pathfinding as a classification task and highlights the trade-offs between model complexity and deployment efficiency.
3.2.6. Approach 6: Novel Pathfinding from Patched Depth Images (Proposed)
@@
To address these challenges, we propose a computationally efficient image processing algorithm for pathfinding. Our method leverages depth images obtained from monocular depth estimation and utilizes a 15×15 patch-based depth representation. Instead of relying on AI models, we implement a pathfinding algorithm to traverse the depth image and identify the longest obstacle-free path, ensuring a clear and safe route for navigation.
@@
Optimal Pathfinding Algorithm:
@@
DEFINE recursive search procedure to explore paths:
@@
IF avg_intensity << threshold OR no valid moves THEN
@@
IF top rows reached OR intensity diff. >> threshold THEN
@@
RECURSE search with updated path
@@
CALL search from (start_row, start_col)
@@
Figure 5. Process for optimal pathfinding calculation illustrating path exploration and optimal route selection.
3.3. Mobile Application Prototype
We created a smartphone application to test our system’s usability in real-world scenarios. The application takes input from the device camera and uses the suggested depth estimation and pathfinding pipeline locally to ensure offline functioning. The interface was built with few buttons and supports multimodal feedback, including speech via the native text-to-speech engine, directional vibration signals, and optional visual overlays illustrated in Figure 7.
To evaluate system responsiveness, the application was tested in common indoor and outdoor environments.
@@ 4. Results and Discussion
Approach 4 (Depth(15x15(patch)) $\rightarrow$ LLM) Phi-3.5-mini-instruct 47.90 18.33 3.257 40.00 32.38 3.257
@@
Approach 6 (Depth(15x15(patch)) $\rightarrow$ Pathfinder Algo)(Proposed) Pathfinder Algorithm 39.81 32.04 0.377 25.78 37.67 0.399
@@
Approach 4 (Depth(15x15(patch)) $\rightarrow$ LLM) Phi-3.5-mini-instruct 40.50 15.33 3.72 42.00 13.33 3.72
@@
Approach 6 (Depth(15x15(patch)) $\rightarrow$ Pathfinder Algo)(Proposed) Pathfinder Algorithm 36.45 34.67 1.136 26.00 38.33 1.332
@@ 4.1. Evaluation Metrics
To evaluate the performance of the proposed directional prediction algorithm, we consider three key metrics: Mean Absolute Error (MAE), Accuracy, and Average Response Time. We evaluated the performance with our annotated dataset (Section 3.2.5). The ground truth directions were manually annotated separately for 300 images from the NYU Depth V2 (Nathan Silberman and Fergus, 2012) dataset (indoor scenes) and 300 images from the DIODE (Vasiljevic et al., 2019) dataset (outdoor scenes).
@@ 4.1.1. Mean Absolute Error (MAE)
MAE measures the average deviation between the predicted navigational direction and the ground truth, both expressed in clock-hour format. For example, if the ground truth direction is annotated as 11:00 (i.e., $330^{\circ}$), and the predicted direction is 10:00 (i.e., $300^{\circ}$), then the angular deviation is $|330^{\circ}-300^{\circ}|=30^{\circ}$, which corresponds to a 1-hour error. Formally, MAE is calculated as:
@@ 4.1.3. Average Response Time
Figure 6. A step-by-step visualization of our proposed method for identifying the optimal path using the Pathfinder algorithm, applied to a sample input from the DIODE (Vasiljevic et al., 2019) dataset.:
@@ 4.2. Indoor Scene Performance
The performance comparison for indoor scenes, as shown in Table 1, reveals key insights into the effectiveness and efficiency of the proposed and baseline approaches. Notably, the proposed Approach 6 (Depth(15x15(patch)) $\rightarrow$ Pathfinder Algo) demonstrates competitive performance across key metrics, balancing mean absolute error (MAE), accuracy, and average response time.
For the indoor scene, Approach 6 achieves an MAE of 25.78 degrees with an accuracy of 37.67% and a notably fast average response time of 0.399 seconds. While the MAE value is slightly better than some LLM-based approaches (e.g., Approach 4 (Depth(5x5(patch)) $\rightarrow$ LLM) with 37.5 degrees MAE using gemini-1.5-pro), the proposed method significantly outperforms others in terms of computational efficiency. The proposed approach achieves an average response time of 0.399 seconds, which is significantly faster than VLM and LLM-based methods, where response times range between 3.094 and 5.75 seconds for other approaches. This indicates that the Pathfinder Algorithm is highly optimized for rapid processing, making it a suitable choice for real-time navigation applications.
Although the classification-based Approach 5 (ViT-Base) exhibits the lowest response time of 0.087 seconds, but its MAE of 52.5 degrees and accuracy of 13.29% indicate a trade-off between speed and prediction quality. This suggests that while ViT-based approaches are computationally efficient, they may lack the precision required for fine-grained obstacle navigation tasks.
@@ 4.3. Outdoor Scene Performance
Table 2 presents the performance comparison for outdoor scenes, where the proposed Approach 6 again demonstrates remarkable efficiency and improved accuracy under specific conditions. For the set of 30 test images, Approach 6 achieves an MAE of only 26.00 degrees with an accuracy of 38.33%, outperforming all other approaches in both metrics. This represents a significant improvement over the best-performing VLM-based approach (Approach 4 with gemini-1.5-pro), which achieved an MAE of 41.00 degrees and an accuracy of 16.67%.
@@
The efficiency of the proposed approach is further highlighted by its response time of 1.332 seconds, which, while slower than the ViT-based method (0.091 seconds), is still markedly faster than all VLM and LLM-based methods, which suggests that the Pathfinder algorithm strikes an effective balance between accuracy and computational efficiency, making it well-suited for real-world applications where rapid decision-making is critical.
@@
Interestingly, the proposed Approach 6 also performs consistently across indoor and outdoor environments, with only a slight increase in response time when transitioning to outdoor scenes. This consistency indicates the robustness of the pathfinding method in diverse environments, which is essential for navigation systems assisting blind and low-vision (BLV) individuals.
Table 3. Impact of Patch Size on Inference Time, MAE, and Accuracy for Indoor (Nathan Silberman and Fergus, 2012) and Outdoor (Vasiljevic et al., 2019) Scenes
Patch Size Indoor Scenes Outdoor Scenes
Avg. Response Time (s) MAE (degree) $\downarrow$ Accuracy (%) $\uparrow$ Avg. Response Time (s) MAE (degree) $\downarrow$ Accuracy (%) $\uparrow$
480 $\times$ 480 0.267 52.80 25.12 0.727 50.95 28.81
240 $\times$ 240 0.294 46.50 28.84 0.831 44.80 30.42
120 $\times$ 120 0.317 44.24 30.50 0.922 41.20 32.40
60 $\times$ 60 0.340 42.95 31.21 0.967 38.37 32.52
30 $\times$ 30 0.362 40.20 34.50 1.012 36.54 34.21
15 $\times$ 15 0.377 39.81 32.04 1.136 36.45 34.67
5 $\times$ 5 0.391 39.84 32.25 1.424 36.48 33.92
Based on the results in Table 3, we adopt a $\times15$ patch size for our depth-based pathfinding framework. Patch size selection is crucial since depth maps are used for object localization, and meaningful detection requires grouping pixels into sufficiently large regions. Very large patches (e.g., $\times480$) result in faster inference but lose granularity, missing smaller or irregularly shaped obstacles. On the other hand, very small patches increase detection granularity but substantially raise inference time.
The $\times15$ patch configuration achieves the best balance across metrics: it yields the lowest MAE for indoor scenes (39.81) and the highest accuracy for outdoor scenes (34.67%), while maintaining response times within practical limits (0.377s indoors and 1.136s outdoors). This trade-off ensures that the system captures diverse object sizes with sufficient accuracy without incurring excessive computational cost, making it well-suited for real-time BLV navigation applications.
4.4. Application Feedback
The system provides navigation guidance through three complementary feedback modes: voice, haptic, and visual. This multimodal design ensures that blind and low-vision (BLV) users can receive clear instructions in diverse contexts, whether indoors, outdoors, quiet, or noisy.
4.4.1. Voice Feedback
Voice instructions are generated using the device’s Text-to-Speech (TTS) engine (Benahmed et al., 2007). Two formats are supported: (i) short prompts such as “move forward”, and (ii) clock-based directions such as “2 o’clock”. These concise formats allow users to quickly interpret orientation and act accordingly. As shown in Fig. 7, audio and clock-direction cues worked best indoors, where quiet and structured spaces allowed participants to follow guidance naturally.
4.4.2. Haptic Feedback
To ensure reliable navigation in noisy environments, the application employs haptic guidance through vibration signals. Distinct vibration patterns and types, corresponding to clock-based directions, are mapped as summarized in Appendix A. Each pattern is unique, making it easy for users to distinguish direction solely through tactile feedback. When participants were outdoors, vibration cues proved more effective in cutting through environmental noise and distractions, while still providing accurate directional assistance.
4.4.3. Visual Feedback
Although unnecessary, we implemented visual feedback. For users with partial vision, where simple visual cues on the phone display, such as arrows (Fig. 7) or clock-hand overlays. This redundant channel reduces uncertainty and supports residual sight usage, making navigation smoother and more intuitive. Visual feedback complements audio and haptic guidance, offering a flexible third layer of support in both indoor and outdoor scenarios.
The strength of the system lies in its multimodal integration. While audio and visual cues are effective in structured indoor spaces, haptic guidance ensures clarity and safety in complex or noisy outdoor environments. Together, the three feedback channels increase flexibility, reduce confusion, and empower BLV users to navigate safely, confidently, and independently.
Figure 7. PathFinder feedback interface showcasing (a) voice feedback, (b) haptic feedback, and (c) visual feedback.